OpenStudy (lizilizi28):

solve factoring 5x^2 +12x-44=0

OpenStudy (campbell_st):

here is my method for a quadratic \[ax^2 + bx + c = 0\] multiply a and c so in your question 5 x -44 = - 220 find the factors that add to 12 22, -10 then write the binomials as (ax + factor 1)(ax + factor 2) -------------------------- = 0 a so you get \[\frac{(5x -10)(5x + 12)}{5} = 0\] remove the common factor from the 1st binomial \[\frac{5(x - 2)(5x + 22)}{5} = 0\] cancel the common factor for the solution.

OpenStudy (johnweldon1993):

@campbell_st wow! I like that method it's alot easier than guess and check

jimthompson5910 (jim_thompson5910):

you can also use the quadratic formula to solve for x, then you can use the zero product property to rewrite as a factorization

OpenStudy (campbell_st):

and it works every time...

jimthompson5910 (jim_thompson5910):

for example, if you solved x^2 + 5x + 6 = 0 for x using the quadratic formula, you would get x = -3 or x = -2 that's equivalent to x+3 = 0 or x+2 = 0 which is the same as (x+3)(x+2) = 0

OpenStudy (campbell_st):

another method is the dicriminant method evaluate the discriminant 1st elps to determine how difficult the problem is so in jims problem its 25 - 24 = 1 then add or subtract b (-5 + 1)/2 and (-5 -1)/2 but for me... I'm happy using the method above...

OpenStudy (lizilizi28):

i need to get two solutions

OpenStudy (campbell_st):

yes so you have the factorised from of the quadratic (x -2)(5x + 12) = 0 so all you need to do is solve x - 2 = 0 and 5x + 12 = 0 hope this helps

OpenStudy (lizilizi28):

so the first one solution is -2 i dont get the second on2

jimthompson5910 (jim_thompson5910):

First subtract 12 from both sides 5x + 12 = 0 5x + 12-12 = 0-12 5x + 0 = -12 5x = -12 what's next?

OpenStudy (lizilizi28):

5/-12 i think

jimthompson5910 (jim_thompson5910):

you have it flipped

jimthompson5910 (jim_thompson5910):

it should be -12/5

jimthompson5910 (jim_thompson5910):

because you would divide both sides by 5